metal occurs as face centered cube and has a density of 7.856 g/cm3. 

Calculate the atomic radius of the metal. Atomic mass of metal is 58 g/mol.


1
Expert's answer
2021-10-29T02:30:23-0400

a= "\frac{4r}{√m}"

7.856= 4×r/√58

7.856= 4r/7.62

4r= 7.62×7.856

4r= 59.86

r=14.97cm